# Contextual Continuum Bandits: Static Versus Dynamic Regret

### v3

@@ 3 Static-to-dynamic regret conversion
Algorithm 1 is a meta-algorithm for static-to-dynamic regret conversion. It uses a partition $\{B_{i}\}_{i=1}^{K^{p}}$ of the hypercube $[0,1]^{p}$, where all the cells $B_{i}$ are hypercubes of equal volume with edges of length $1/K$. Fix a consistent randomized procedure $\pi$ for the class of functions $\mathcal{F}$. At each round $t$, if the context $\boldsymbol{c}_{t}$ falls into the cell $B_{i}$, Algorithm 1 chooses the query point $\boldsymbol{z}_{t}$ by the procedure $\pi$ based only on the past contexts from the same cell $B_{i}$ and the corresponding past query points.
@@
where $\Delta_{i_{j}}=f(\boldsymbol{x}_{i}^{*},\boldsymbol{c}_{i_{j}})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})$. Notice that $\Delta_{i_{j}}\leq 2\max_{\boldsymbol{z}\in\Theta}|f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{z},\boldsymbol{b}_{i})|$, so that under Assumption 1 we have $\Delta_{i_{j}}\leq 2L(\sqrt{p}/K)^{\gamma}$. Thus, we can control the dynamic regret by the sum of the static regret, which grows with $K$, and the bias term comprised of $\Delta_{i_{j}}$’s, which decreases with $K$. There exists an optimal $K$ exhibiting a trade-off between the static regret and the bias. In Theorem 1, we show that if $K$ is tuned optimally,
@@
Note that the Lipschitz constant $L$ measures how sensitive the function $f$ is with respect to variations of the contexts. If $L=0$ the static and dynamic regrets are equivalent, and it is not surprising that (7) and (1) have the same rates. We also note that a bound that depends in the same way as (1) on the parameters $p$, $d$, and $T$ can be achieved with a choice of $K$ independent of $L$. In Theorem 1 we assume that the learner knows the value of $L$ but this is only done to highlight the gap between the notions of static and dynamic regret.
The approach to contextual bandits based on partitioning the space of contexts and running some algorithms $\pi$ separately within each cell is not new, see [27, 13, 42]. These papers, as well as the follow-up works developing the same approach, focused on a specific algorithm $\pi$, a specific class $\mathcal{F}$ and provided a proof using their particular properties. They did not invoke conversion from static to dynamic regret. In contrast, Theorem 1 highlights a general static-to-dynamic conversion phenomenon and provides a ready-to-use tool for exploring particular examples. In the next section, we demonstrate its application to the important case that $\mathcal{F}^{\prime}$ is the class of strongly convex and smooth functions.
@@ 4 Strongly convex and smooth functions
In this section, we apply the static-to-dynamic regret conversion developed in Section 3 to the contextual bandits problem with the objective functions $f$ such that $f(\cdot,\boldsymbol{c})$ is $\beta$-smooth and $\alpha$-strongly convex for any $\boldsymbol{c}\in[0,1]^{p}$.
Definition 2.
Let $\beta>0$. Function $F:\mathbb{R}^{d}\to\mathbb{R}$ is called $\beta$-smooth, if it is continuously differentiable and for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$ it satisfies
$\left\lVert\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\rVert\leq\beta\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert.$
Definition 3.
Let $\alpha>0$. A differentiable function $F:\mathbb{R}^{d}\to\mathbb{R}$ is called $\alpha$-strongly convex, if for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$ it satisfies
$F(\boldsymbol{x})\geq F(\boldsymbol{y})+\langle\nabla F(\boldsymbol{y}),\boldsymbol{x}-\boldsymbol{y}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert^{2}.$
@@
Assumption 2.
For $(\alpha,\beta,M)\in(0,\infty)\times[\alpha,\infty)\times(0,\infty)$, we assume that $f\in\mathcal{F}_{\alpha,\beta}(M)$, where $\mathcal{F}_{\alpha,\beta}(M)$ denotes the class of all functions $f:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ such that for all $\boldsymbol{c}\in[0,1]^{p}$ the following holds.
@@
Following the approach of Section 3, we first construct a consistent randomized procedure, that is, a procedure with sub-linear static regret, cf. (6). To this end, our starting point is the BCO algorithm introduced in [20] to solve strongly convex bandit problems. The BCO algorithm makes use of the interior point method, which is a vital tool in convex optimization, see, e.g, [36]. A pioneering role of introducing this method to the field of online learning is credited to [1]. We cannot directly apply the BCO algorithm as defined in [20] since [20] deals with the setting, where the observations $y_{t}$ are noise-free. Therefore, we propose a modification of the BCO algorithm adapted to noisy observations (3) with sub-Gaussian noise $\xi_{t}$. It is defined in Algorithm 2. In what follows, $[T]$ denotes the set of all positive integers less than or equal to $T$. We denote the unit Euclidean ball in $\mathbb{R}^{d}$ by $B^{d}$ and the unit Euclidean sphere by $\partial B^{d}$.
Assumption 3.
There is $\sigma>0$ such that for all $t\in[T]$ the following holds: (i) $\xi_{t}$ is $\sigma$-sub-Gaussian, i.e., for any $\lambda>0$ we have $\mathbf{E}\left[\exp\left(\lambda\xi_{t}\right)\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)$; (ii) $\xi_{t}$ is independent of $\boldsymbol{\zeta}_{t}$.
@@
    where for $m\in[3]$, and $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{m}\in\mathbb{R}^{d}$
    $\nabla^{m}\mathcal{R}(\boldsymbol{y})[\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{m}]=\frac{\partial^{m}}{\partial t_{1}\dots\partial t_{m}}\mathcal{R}\left(\boldsymbol{y}+t_{1}\boldsymbol{x}_{1}+\dots+t_{m}\boldsymbol{x}_{m}\right)|_{t_{1}=0,\dots,t_{m}=0}.$
@@
Since $\Theta$ is bounded, $\nabla^{2}\mathcal{R}(\boldsymbol{x})$ is positive definite and the minimizer of $\mathcal{R}$ in $\text{int}(\Theta)$ is unique, see [35, Chapter 3, Section V]. Moreover, without loss of generality, we assume that $\min_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})=0$. Self-concordant barriers can be chosen depending on the geometry of $\Theta$. For the examples of $\mathcal{R}$ and the corresponding values of $\mu$, we refer the reader to Appendix D.
@@
       Generate $\boldsymbol{\zeta}_{t}$ from $\partial B^{d}$
@@
      $\boldsymbol{x}_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\Theta}\sum_{k=1}^{t}\eta_{k}\left(\langle\boldsymbol{g}_{k},\boldsymbol{x}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{k-1}\right\rVert^{2}\right)+\mathcal{R}(\boldsymbol{x})$
@@
The discrepancy between Algorithm 2 and the BCO algorithm of [20] is in the assigned step size. In [20] the step size is constant while it varies with $t$ in Algorithm 2. Next, in Algorithm 2, the step size is downscaled by $q_{T}=M+2\sigma\sqrt{\log(T+1)}$ whereas in [20] it is downscaled by $M$ without any additional logarithmic term. This difference is due to the fact that the observations in (3) are noisy whereas in [20] they are assumed to be noise-free. In our analysis, we rely on the fact that $\eta_{t}\left\lVert P_{t}\boldsymbol{g}_{t}\right\rVert$ is uniformly bounded with high probability. To ensure this, it is enough to scale down the step size by adding the above logarithmic term to $M$. Indeed, due to Assumption 3, we have that $\max_{t\in[T]}|y_{t}|\leq q_{T}$ with a suitably high probability.
Algorithm 2 algorithm selects $\boldsymbol{z}_{t}\in\{\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}|\boldsymbol{\zeta}\in\partial B^{d}\}\subseteq D\triangleq\{\boldsymbol{z}|(\boldsymbol{z}-\boldsymbol{x}_{t-1})^{\top}\nabla^{2}\mathcal{R}(\boldsymbol{x}_{t-1})(\boldsymbol{z}-\boldsymbol{x}_{t-1})\leq 1\}$, in which $D$ is the Dikin ellipsoid. Consequently, $D\subseteq\Theta$; see [35, Proposition 14.2.2] for a proof. Therefore, $\boldsymbol{z}_{t}\in\Theta$ for all $t\in[T]$.
Theorem 2.

### v6

@@ 3 A general conversion theorem
Algorithm 1 is a meta-algorithm that assures regret conversion. It uses a partition $\{B_{i}\}_{i=1}^{K^{p}}$ of the hypercube $[0,1]^{p}$, where all the cells $B_{i}$ are hypercubes of equal volume with edges of length $1/K$. Fix a consistent randomized procedure $\pi$ for the class of functions $\mathcal{F}$. At each round $t$, if the context $\boldsymbol{c}_{t}$ falls into the cell $B_{i}$, Algorithm 1 chooses the query point $\boldsymbol{z}_{t}$ by the procedure $\pi$ based only on the past contexts from the same cell $B_{i}$ and the corresponding past query points. Since no more than $T$ cells $B_{i}$ can be visited and need to be memorized Algorithm 1 is polynomial time provided the input algorithm $\pi$ is polynomial time.
@@
where $\Delta_{i_{j}}=f(\boldsymbol{x}_{i}^{*},\boldsymbol{c}_{i_{j}})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})$. One may notice that $\Delta_{i_{j}}\leq 2\max_{\boldsymbol{z}\in\Theta}|f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{z},\boldsymbol{b}_{i})|$ (see the proof of Theorem 2), so that under Assumption 1 we have $\Delta_{i_{j}}\leq 2L(\sqrt{p}/K)^{\gamma}$. Thus, we can control the dynamic regret by the sum of the static regret, which grows with $K$, and the bias term comprised of $\Delta_{i_{j}}$’s, which decreases with $K$. There exists an optimal $K$ exhibiting a trade-off between the static regret and the bias. In Theorem 2, we show that if $K$ is tuned optimally,
@@
Note that the Lipschitz constant $L$ measures how sensitive the function $f$ is with respect to variations of the contexts. If $L=0$ the static and contextual regrets are equivalent, and it is not surprising that (8) and (2) have the same rates. We also note that a bound that depends in the same way as (2) on the parameters $p$, $d$, and $T$ can be achieved with a choice of $K$ independent of $L$. In deriving (2) we assumed that the learner knows the value of $L$ but this is only done to illustrate the above remark.
The approach to contextual bandits based on discretizing the space of contexts and running some algorithms $\pi$ separately for each discretization item is quite common, see Hazan and Megiddo (2007); Kleinberg et al. (2008); Bubeck et al. (2008); Slivkins (2014); Cesa-Bianchi et al. (2017); Li et al. (2019). With the exception of Slivkins (2014), each of these papers, as well as follow-up works based on similar arguments, explored a specific algorithm $\pi$, a specific class $\mathcal{F}$, and provided bounds on the contextual regret (or on $L$-regret) using their particular properties. More generally, (Slivkins, 2014, Theorem 21) provides a static-to-contextual regret conversion device, where the class $\mathcal{F}$ is still quite specific (defined by conditions (S1) and (S2) below) but the input algorithm $\pi$ is not fixed and allowed to belong to the class of algorithms with $\tilde{O}(T^{\frac{1}{p+2}})$ static regret. The bound on the contextual regret in (Slivkins, 2014, Theorem 21) includes the doubling dimension, which is exponential in the dimension of the context. In contrast, Theorem 2 is a conversion result valid for a general class $\mathcal{F}$, a general input algorithm $\pi$, and its bound is free from exponential factors. It yields a ready-to-use tool for exploring various particular settings. Below we demonstrate its application to the three major contextual settings, where the class $\mathcal{F}$ is defined by Assumption 1 combined with one of the following assumptions on the class of functions $f(\cdot,\boldsymbol{c})$ for any fixed $\boldsymbol{c}$:
  • (a)
    $\mathcal{F}^{\prime}$ is a class of Lipschitz functions,
  • (b)
    $\mathcal{F}^{\prime}$ is the class of all convex functions,
  • (c)
    $\mathcal{F}^{\prime}$ is a class of strongly convex and smooth functions.
For the sake of comparison, we recall that the class of loss functions $\mathcal{F}$ considered in (Slivkins, 2014, Theorem 21) is defined by the following two conditions:
00 (S1) $f(\boldsymbol{z},\boldsymbol{c})$ is Lipschitz in $(\boldsymbol{z},\boldsymbol{c})$ with Lipschitz constant 1,
00 (S2) for each context $\boldsymbol{c}$, there exists $\boldsymbol{z}^{*}(\boldsymbol{c})\in\operatorname*{arg\,min}_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c})$ such that the map $\boldsymbol{c}\mapsto f(\boldsymbol{z}^{*}(\boldsymbol{c}),\boldsymbol{c})$ is Lipschitz, with Lipschitz constant 1.
0 Condition (S2) does not follow from (S1) and requires some strong additional properties of $f$ that are not detailed in Slivkins (2014). Moreover, as emphasized in Slivkins (2014) the assumption that the Lipschitz constant is 1 can be relaxed but in that case the algorithm should know the Lipschitz constant.
For the rest of this section, we focus on the implications of Theorem 2 for two families of functions corresponding to items (a) and (b) above. Observe that considering (a) contains as a special case the class $\mathcal{F}$ of functions $f(\boldsymbol{z},\boldsymbol{c})$ that are Lipschitz in $(\boldsymbol{z},\boldsymbol{c})$ with no further restriction. We do not need the hardly interpretable additional condition (S2), nor the knowledge of the Lipschitz constant.
(a) Lipschitz bandits. Lipschitz non-contextual bandits have been studied extensively in the literature, cf. Lattimore and Szepesvári (2020) and the references therein. The main idea is to reduce the Lipschitz bandit problem to a multi-armed bandit problem with finite number of arms, often employing zooming techniques with dyadic splits; see, for example, Kleinberg (2004); Kleinberg et al. (2008); Bubeck et al. (2008). While these methods benefit from near optimal in $T$ static regret (of the order $\tilde{O}(T^{\frac{d+1}{d+2}})$), the regret bounds depend exponentially on the dimension $d$. A recent improvement due to Podimata and Slivkins (2021) shows that polynomial dependence on the dimension is achievable. In Podimata and Slivkins (2021), the authors proposed an algorithm with static regret that scales as $d^{\frac{1}{2}}T^{\frac{d+1}{d+2}}\log^{5}(T)$ for $\Theta=[0,1]^{d}$ (cf. (Podimata and Slivkins, 2021, Theorem 2) with $z=d$ for the worst-case scenario). Taking their method as an input $\pi$ of our meta-algorithm and combining the bound of (Podimata and Slivkins, 2021, Theorem 2) with Theorem 2 we obtain the following corollary.
Corollary 3
Assume that $\Theta=[0,1]^{d}$ and that the observations in (2) are noise-free, that is, $\xi_{t}=0$ for all positive integers $t$. Let $\mathcal{F}$ be the set of functions such that $|f(\boldsymbol{x},\boldsymbol{c})-f(\boldsymbol{x}^{\prime},\boldsymbol{c})|\leq L^{\prime}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in[0,1]^{d}$, $\boldsymbol{c}\in[0,1]^{p}$, where $L^{\prime}>0$ is a constant. Let $\boldsymbol{z}_{t}$’s be the updates of Algorithm 1 with the input algorithm $\pi$ taken as the Lipschitz bandit algorithm from Podimata and Slivkins (2021). Then, under the optimal choice of $K$ we have
$)\displaystyle\sup_{f\in\mathcal{F}\cap\mathcal{F}_{\gamma}(L)}\mathbf{E}\big{[}\sum_{t=1}^{T}\big{(}f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})$$\big{)}\big{]}\leq C\max\bigg{(}d^{1/2}T^{(d+1)/(d+2)}L_{1},$
$\big{(}p^{\gamma p/2}\big{(}d^{\frac{1}{2}}L_{1}\big{)}^{\gamma(d+2)}T^{p+(d+1)\gamma}\big{)}^{1/(p+(d+2)\gamma)}\bigg{)},$
where $L_{1}=\log^{5}(T+1)$, and $C>0$ is a constant that does not depend on $T$, $d$, and $p$.
Proof  By (Podimata and Slivkins, 2021, Theorem 2) the input algorithm $\pi$ satisfies the static regret bound (7) with
$F(T)\leq C_{1}d^{1/2}T^{(d+1)/(d+2)},\quad\text{and}\quad F_{1}(T+1)\leq\log^{5}(T+1),$
where $C_{1}>0$ is a constant independent of $T$, $d$, and $p$. Thus, we can apply Theorem 2 with $\tau_{1}=1/2$, $\tau_{2}=(d+1)/(d+2)$, and $\tau_{3}=5$, which yields the result.
In particular, in the case $\gamma=1$, that is, when $f$ is Lipschitz continuous in both arguments, Corollary 3 establishes the rate in $T$ as $T^{(p+d+1)/(p+d+2)}$ up to a logarithmic factor and a weak factor depending on $p,d$, which does not exceed $O(\sqrt{pd})$. On the other hand, for $\gamma=1$, Slivkins (2014) provides a matching lower bound on the contextual regret of the same order $T^{(p+d+1)/(p+d+2)}$. Together with Corollary 3, it implies that the rate $T^{(p+d+1)/(p+d+2)}$ is minimax optimal as function of $T$, up to logarithmic factors, and our algorithm attains the minimax rate. Notice that the minimax optimality of this rate on the class of functions $f$ that are Lipschitz continuous in both arguments was not established in Slivkins (2014), contrary to what is sometimes claimed. Indeed, Slivkins (2014) provided an upper bound on a substantially smaller class where additionally condition (S2) holds. Furthermore, the upper bound in (Slivkins, 2014, Theorem 21) is proportional to doubling dimension, which scales exponentially in $p$, while Corollary 3 involves only a weak dimension dependent factor that never exceeds $O(\sqrt{pd})$.
(b) Convex bandits. The study of non-contextual bandit convex optimization was initiated by Flaxman et al. (2004) and Kleinberg (2004). There exists an extensive literature on this problem for both stochastic and adversarial settings, see Orabona (2019); Lattimore and Szepesvári (2020); Lattimore (2024) and the references therein. Bounds on the static regret in adversarial setting were developed in Bubeck et al. (2017); Bubeck and Eldan (2018); Lattimore and Szepesvári (2020); Fokkema et al. (2024). Papers Bubeck et al. (2017); Bubeck and Eldan (2018); Lattimore and Szepesvári (2020) considered the setting with no noise ($\xi_{t}=0$ for all $t$) providing successive improvements of static regret bounds. The best known result Lattimore and Szepesvári (2020) yields a bound of the order $d^{2.5}\sqrt{T}$ derived as an evaluation of the minimax regret without proposing an explicit algorithm (here and below we ignore the logarithmic factors). An important progress was recently achieved in Fokkema et al. (2024), where the authors proposed a polynomial time algorithm attaining a regret of the order $d^{3.5}\sqrt{T}$ in adversarial setting with sub-Gaussian noise, cf. (Fokkema et al., 2024, Theorem 1). Combining this result, which is the state of the art, with Theorem 2 we obtain Corollary 4 below for the contextual convex bandit setting. As (Fokkema et al., 2024, Theorem 1) does not explicitly detail the dependence on logarithmic factors, we adopt the same convention and present the bound up to logarithmic factors.
Corollary 4
Let $\mathcal{F}$ be the set of functions $f$ such that $f(\cdot,\boldsymbol{c})$ is convex for all $\boldsymbol{c}\in[0,1]^{p}$. Assume that, conditionally on the past observations, $\xi_{t}$ in (2) is a zero-mean $\sigma$-sub-Gaussian random variable, that is, for any $\lambda\in{\mathbb{R}}$ and some $\sigma>0$ we have $\mathbf{E}\left[\exp\left(\lambda\xi_{t}\right)|\{\boldsymbol{c}_{k},\boldsymbol{z}_{k},y_{k}\}_{k=1}^{t-1},\boldsymbol{c}_{t}\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)$ [^1] Requiring the inequality for all $\lambda\in{\mathbb{R}}$ grants that $\xi_{t}$ has zero conditional mean. for all positive integers $t$. Let $\boldsymbol{z}_{t}$’s be the updates of Algorithm 1 with the input algorithm $\pi$ taken as the convex bandit algorithm from Fokkema et al. (2024). Then, under the optimal choice of $K$ we have
$)\displaystyle\sup_{f\in\mathcal{F}\cap\mathcal{F}_{\gamma}(L)}\mathbf{E}\big{[}\sum_{t=1}^{T}\big{(}f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})$$\big{)}\big{]}\leq C\max\left(d^{7/2}\sqrt{T},p^{\gamma p/(2p+4\gamma)}d^{7\gamma/\left(p+2\gamma\right)}T^{\frac{p+\gamma}{p+2\gamma}}\right),$
where $C>0$ is a factor that depends polynomially on $\log(T)$ and $\log(d)$ and does not depend on $p$.
Proof  By (Fokkema et al., 2024, Theorem 1) the input algorithm $\pi$ satisfies the static regret bound (7) with $F(T)\leq C^{\prime}d^{7/2}\sqrt{T}$ and $F_{1}(T)=\log^{a}(T)$ for some $a>0$ and a constant $C^{\prime}>0$ that depends polynomially on $\log(d)$ and does not depend on $p$. Thus, we can apply Theorem 2 with $\tau_{1}=7/2$ and $\tau_{2}=1/2$. Taking $K$ as defined in (9) with $\tau_{1}=7/2$, $\tau_{2}=1/2$, and $\tau_{3}=a$ we obtain the result of the theorem (with the dependency with respect to logarithmic factors hidden in $C$).
Thus, for convex contextual bandits we have an algorithm with the contextual regret that scales in $T$ as $T^{\frac{p+\gamma}{p+2\gamma}}$ up to a logarithmic factor. The dependence on $d$ and $p$ is mild. It is given by a factor that never exceeds $p^{1/2}d^{7/3}$, to within a power of $\log(d)$. Combining this result with the lower bound of Theorem 11 (see Section 5 below) we conclude that the minimax optimal rate in $T$ for convex contextual bandits is of the order $T^{\frac{p+\gamma}{p+2\gamma}}$ up to a logarithmic factor.
@@ 4 Strongly convex and smooth functions
In this section, we apply the regret conversion approach developed in Section 3 to the contextual bandits problem with the objective functions $f$ such that $f(\cdot,\boldsymbol{c})$ is $\beta$-smooth and $\alpha$-strongly convex for any $\boldsymbol{c}\in[0,1]^{p}$.
Definition 5
Let $\beta>0$. Function $g:\mathbb{R}^{d}\to\mathbb{R}$ is called $\beta$-smooth, if it is continuously differentiable and for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$ it satisfies
$\left\lVert\nabla g(\boldsymbol{x})-\nabla g(\boldsymbol{y})\right\rVert\leq\beta\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert.$
Definition 6
Let $\alpha>0$. A differentiable function $g:\mathbb{R}^{d}\to\mathbb{R}$ is called $\alpha$-strongly convex, if for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$ it satisfies
$g(\boldsymbol{x})\geq g(\boldsymbol{y})+\langle\nabla g(\boldsymbol{y}),\boldsymbol{x}-\boldsymbol{y}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert^{2}.$
@@
Assumption 2
For $(\alpha,\beta,M)\in(0,\infty)\times[\alpha,\infty)\times(0,\infty)$, we assume that $f\in\mathcal{F}_{\alpha,\beta}(M)$, where $\mathcal{F}_{\alpha,\beta}(M)$ denotes the class of all functions $f:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ such that for all $\boldsymbol{c}\in[0,1]^{p}$ the following holds:
@@
The non-contextual bandit setting with strongly convex and smooth functions has been investigated in Hazan and Levy (2014) and Ito (2020). Both Hazan and Levy (2014) and Ito (2020) studied the problem with no noise ($\xi_{t}\equiv 0$ for all $t$). In Hazan and Levy (2014), the authors proposed an algorithm leveraging an interior point method and a self-concordant function. They achieved a static regret, which is nearly optimal, ranging from $d\sqrt{T}$ to $d^{3/2}\sqrt{T}$ depending on the geometry of the constraint set. Assuming that all minimizers of the objective functions are in the interior of the constraint set, Ito (2020) introduced a procedure that achieves static regret of the optimal order $d\sqrt{T}$ (however, with no polynomial time guarantee). In this section, we extend the algorithm of Hazan and Levy (2014) to the setting with noisy function evaluation as in (2) and derive a bound on the static regret of the proposed algorithm. Then, using it as an input algorithm $\pi$ of our meta-algorithm we obtain a bound on the contextual regret of the resulting procedure.
Following the approach of Section 3, we first construct a consistent randomized procedure, that is, a procedure with sub-linear static regret, cf. (7). To this end, our starting point is the BCO algorithm introduced in Hazan and Levy (2014). The BCO algorithm makes use of the interior point method, which is a vital tool in convex optimization, see, e.g, Nesterov and Nemirovskii (1994). A pioneering role of introducing this method to the field of online learning is credited to Abernethy et al. (2008). We cannot directly apply the BCO algorithm as defined in Hazan and Levy (2014) since Hazan and Levy (2014) deals with the setting, where the observations $y_{t}$ are noise-free. Therefore, we propose a modification of the BCO algorithm adapted to noisy observations (2) with sub-Gaussian noise $\xi_{t}$. It is defined in Algorithm 2. In what follows, $[T]$ denotes the set of all positive integers less than or equal to $T$. We denote the unit Euclidean ball in $\mathbb{R}^{d}$ by $B^{d}$ and the unit Euclidean sphere by $\partial B^{d}$.
Assumption 3
There is $\sigma>0$ such that for all $t\in[T]$ the following holds: (i) $\xi_{t}$ is $\sigma$-sub-Gaussian, i.e., for any $\lambda>0$ we have $\mathbf{E}\left[\exp\left(\lambda\xi\right)\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)$; (ii) $\xi_{t}$ is independent of $\boldsymbol{\zeta}_{t}$.
@@
    where for $m\in{1,2,3}$, and $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}\in\mathbb{R}^{d}$
    $\nabla^{m}\mathcal{R}(\boldsymbol{y})[\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}]=\frac{\partial^{m}}{\partial t_{1}\dots\partial t_{m}}\mathcal{R}\left(\boldsymbol{y}+t_{1}\boldsymbol{v}_{1}+\dots+t_{m}\boldsymbol{v}_{m}\right)|_{t_{1}=0,\dots,t_{m}=0}.$
@@
Since $\Theta$ is bounded, $\nabla^{2}\mathcal{R}(\boldsymbol{x})$ is positive definite, the minimizer of $\mathcal{R}$ in $\text{int}(\Theta)$ is unique, see (Nemirovski, 2004, Chapter 3, Section V), and $\boldsymbol{x}_{t}$ in Algorithm 2 is uniquely defined as well. Moreover, without loss of generality, we assume that $\min_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})=0$. Self-concordant barriers can be chosen depending on the geometry of $\Theta$. For the examples of $\mathcal{R}$ and the corresponding values of $\mu$, we refer the reader to Appendix D.
@@
       Generate $\boldsymbol{\zeta}_{t}$ uniformly distributed on $\partial B^{d}$
@@
      $\boldsymbol{x}_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\Theta}\eta_{t}\sum_{k=1}^{t}\left(\langle\boldsymbol{g}_{k},\boldsymbol{x}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{k-1}\right\rVert^{2}\right)+\mathcal{R}(\boldsymbol{x})$
@@
The difference between Algorithm 2 and the BCO algorithm of Hazan and Levy (2014) is in the assigned step size. In Hazan and Levy (2014) the step size is constant and depends on the horizon $T$ while it varies with $t$ in Algorithm 2. Next, in Algorithm 2, the step size is downscaled by $q_{T}=M+2\sigma\sqrt{\log(T+1)}$ whereas in Hazan and Levy (2014) it is downscaled by $M$ without any additional logarithmic term. This difference is due to the fact that the observations in (2) are noisy whereas in Hazan and Levy (2014) they are assumed to be noise-free. In our analysis, we rely on the fact that $\eta_{t}\left\lVert P_{t}\boldsymbol{g}_{t}\right\rVert$ is uniformly bounded with high probability. To ensure this, it is enough to scale down the step size by adding the above logarithmic term to $M$. Indeed, due to Assumption 3, we have that $\max_{t\in[T]}|y_{t}|\leq q_{T}$ with a suitably high probability (see the details in the proof of Theorem 8).
Algorithm 2 algorithm selects $\boldsymbol{z}_{t}\in\{\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}|\boldsymbol{\zeta}\in\partial B^{d}\}\subseteq D\triangleq\{\boldsymbol{z}|(\boldsymbol{z}-\boldsymbol{x}_{t-1})^{\top}\nabla^{2}\mathcal{R}(\boldsymbol{x}_{t-1})(\boldsymbol{z}-\boldsymbol{x}_{t-1})\leq 1\}$, in which $D$ is the Dikin ellipsoid. Consequently, $D\subseteq\Theta$; see (Nemirovski, 2004, Proposition 14.2.2) for a proof. Therefore, $\boldsymbol{z}_{t}\in\Theta$ for all $t\in[T]$.
Theorem 8